Tons of crude oil?

A supersonic aircraft consumes 5320 imperial gallons of kerosene per hour of flight and flies an average of 14 hours per day. It takes roughly seven tons of crude oilto produce one ton of kerosene. The density of kerosene is 0.965g/cm cube. How many planes would it take to consume the entire annual world production of 4.02 x 10 ^9metric tons of crude oil?
Answer #1

4.02*109metric tons of crude oil

====> 4.02*109/7 tons of kerosene

====> (4.02*109/7)/0.965 m3of kerosene

====> (4.02*109/7)/0.965 * 264.17gallons of kerosene

1 plane consumes 5320*14 gallons of kerosene

so number of plane = [(4.02*109/7)/0.965 * 264.17]/(5320*14) = 2.11*106

answered by: Kodee
Answer #2

(4.02x109metric tons)*(2204.622lbs/1 metric ton)*(453.59g/1lb)*(1L/1000cm3)*(1cm3/.965g)=4.165780168x1012L

(5320 imperial gal/plane hr)*(14hr)*(1.201gal/1 imperial gal)*(1L/.26417gal)=3.386095317x105L/Plane

4.165780168x1012L/3.386095317x105L/Plane=1.230260751x107planes traveling at supersonic speed for 14hrs(1day)

answered by: matt ferrisi
Answer #3

4.02*109 metric tons of crude oil
====> 4.02*109/7 tons of kerosene
====> (4.02*109/7)/0.965 m3 of kerosene

====> (4.02*109/7)/0.965 * 1000 litersof kerosene

====> (4.02*109/7)/0.965 * 1000/4.546 imperial gallons of kerosene

1 plane consumes 5320*14 gallons of kerosene
so number of plane = [(4.02*109/7)/0.965 * 1000/4.546]/(5320*14) = 1.76*106

answered by: Kurt Kincaid